OpenStudy (anonymous):

suppose that the number of cars manufactured at an automobile plant varies jointly as the number of workers and the time they work. If 340 workers can produce 204 cars in 4 hours how many workers would it take to produce 180 cars in 3 hours?

OpenStudy (anonymous):

let c=knw thus by the therem of joint variation we have 204/180 =(340/w )(4/3) w=180/(204) *(4/3) *340=400